# Graph Neural Networks for Binary Programming

### v1

@@ Appendix B Metrics
Solution Accuracy (%). The solution accuracy measured the number of correctly labeled nodes in each example. Mathematically, given a graph $G=(V,E)$, with $n$ nodes, predicted solution ${\bf x}_{p}$ and ground-truth solution ${\bf x}_{o}$, the accuracy is defined as follows:
${\rm{Accuracy}}=\frac{\sum_{i=0}^{n-1}\mathbbm{1}_{{\bf x}_{o_{i}}={\bf x}_{p_{i}}}}{n},$ (27)
where ${\bf x}_{o_{i}}$ denotes the $i$-th entry of ${\bf x}_{o}$, and similarly for ${\bf x}_{p_{i}}$. $\mathbbm{1}_{{\bf x}_{o_{i}}={\bf x}_{p_{i}}}$ is an indicator function that reads 1 if the equality between the entries holds, and 0 otherwise.
@@
Relative QUBO Objective. Recall the QUBO objective from Equation 1. At the optimal solution ${\bf x}_{o}$, its value reads:
@@
Inference time. We measure the time it takes for each approach (exhaustive search, classical solvers, other GNNs, and ours) to solve a QUBO problem. The time is measured in milliseconds and averaged over the solution of 100 problems with the same size per dataset. The measurements were conducted using an Nvidia RTX-4090 GPU.
Appendix C Datasets
@@
We create 3 datasets: small ($k=10$), medium ($k=256$), and large ($k=1024$). For the small dataset, we generate exact solutions using an exhaustive search, with 1 million examples. For the small and medium examples, we generate 1000 examples using the procedure described in Algorithm 1. Since we consider problems where ${\bf A}$ is fixed but the observed-vector ${\bf b}$ is variable, we repeat the process of generating datasets 10 times, so that we have examples on several randomly generated matrices ${\bf A}$. The results in Section 7 are averaged across the 10 datasets, following the common experimental settings in Xu et al. (2019). In all cases, we employ a random training/validation split of 80%/20%, respectively.
@@
For the medium and large datasets, we chose ${\bf A}$ to be the matrix that discretizes the 2D Laplacian on a lattice. Such matrices often appear in thermodynamics calculations (Evans, 1998). The data is generated using algorithm 1. In order to make the problem non-trivial, we choose $\sigma=4$ which makes roughly $5\%$ of the entries in ${\bf x}_{o}$ change. We have used $1000$ Tabu search steps in step (4) of the algorithm to approximate ${\bf x}$.
Appendix D Hyperparameters
@@
In our experiments we consider the following hyperparameters for all GNN methods, including ours: ${\rm{learning\textunderscore rate}}\in\{1e-5,1e-4,1e-3\}$, ${\rm{weight}\textunderscore decay}\in\{1e-5,1e-4,0\}$, ${\rm{dropout}}\in\{0,0.1,0.5\}$. We train each network for up to 200 epochs. Dropout layers, if used, are applied after each hidden layer. We used the Adam (Kingma & Ba, 2014) optimizer in all experiments.
@@
Appendix E Additional Results
E.1 Standard Deviation of the Relative QUBO Loss
@@
Method $\downarrow$ / Problem Size $\rightarrow$ Small ($k=10$) Medium ($k=256$) Large $(k=1024)$
@@
BPGNN $2.83\cdot 10^{-3}\pm 4.48\cdot 10^{-4}$ $5.31\cdot 10^{-3}\pm 5.92\cdot 10^{-4}$ $1.26\cdot 10^{-2}\pm 5.47\cdot 10^{-3}$
BPGNN + TS $1.37\cdot 10^{-5}\pm 2.77\cdot 10^{-6}$ $1.21\cdot 10^{-5}\pm 3.81\cdot 10^{-6}$ $1.10\cdot 10^{-4}\pm 7.09\cdot 10^{-6}$
Table 3: The Relative QUBO loss (lower is better) average and standard deviations over 10 experiments, using various naive exhaustive search methods, classical methods, GNNs, and our BPGNN.
E.2 The Importance of QUBO-Aware Features
Figure 3: The relative QUBO loss convergence in training and validation of BPGNN on the small problem ($k=10)$, with and without the QUBO-aware node features from Equation 14.
We now provide additional results on the importance of including the QUBO-aware features, on small and medium-sized problems, in Tables 4 and 5, respectively. As discussed in Section 7.2, the results indicate that including the QUBO-aware features improves accuracy and the relative QUBO objective.
Furthermore, as we show in Figure 3, a beneficial property of the QUBO-aware features is their ability to accelerate the training convergence of the network.
@@ E.2 The Importance of QUBO-Aware Features
Table 4: The impact of adding the QUBO aware features from Equation 14 to the node features on small sized problems ($k=10$).
@@
Table 5: The impact of adding the QUBO aware features from Equation 14 to the node features on medium sized problems ($k=256$).

### v2

@@ 7 Numerical Experiments
    Solution Accuracy (%). Accuracy is the fraction of correctly labeled nodes per example. Given a graph $G=(V,E)$ with $n$ nodes, predicted solution ${\bf x}_{p}$, and ground truth ${\bf x}_{o}$, it is defined as
    ${\rm Accuracy}=\frac{\sum_{i=0}^{n-1}\mathbf{1}_{{\bf x}_{o_{i}}={\bf x}_{p_{i}}}}{n},$ (26)
    where ${\bf x}_{o_{i}}$ and ${\bf x}_{p_{i}}$ are the $i$-th entries of ${\bf x}_{o}$ and ${\bf x}_{p}$, and $\mathbf{1}_{{\bf x}_{o_{i}}={\bf x}_{p_{i}}}$ is 1 if equal and 0 otherwise. As discussed in Section 6, the data-generation algorithm ensures, with very high probability [9], that ${\bf x}_{o}^{(j)}$ is the minimizer of the QUBO problem defined by $({\bf A},{\bf b}^{(j)})$. Hence, measuring accuracy in Section 7 on this dataset is reliable with high probability.
  2. 2.
@@
    Relative QUBO Objective. Recall the QUBO objective from Equation˜1. At the optimal solution ${\bf x}_{o}$, its value reads:
@@
  3. 3.
    Inference time. We measure the time (ms) each approach (exhaustive search, classical solvers, other GNNs, and ours) takes to solve a QUBO, averaged over 100 problems of equal size per dataset on an Nvidia RTX-4090 GPU. For GNNs, unlike classical solvers [46], and following our approach of learning from QUBO examples to generalize solutions rather than solving from scratch, inference time excludes training, as standard in supervised neural networks. Thus, during inference, the network performs no additional optimization, but directly maps input $({\bf A},{\bf b})$ to an approximate solution $\bf x$}.
Datasets.
Our experiments include three datasets generated by our procedure described in Section 6, in varying sizes - small ($k=10$), medium ($k=256$) and large ($k=1024$). Note that the settings of $k=\{256,1024\}$ are considered highly challenging, as there are $2^{\{256,1014\}}$ possible solutions to the QUBO problem. We provide further details on the datasets in Appendix A.
Training and Inference.
Our training procedure utilizes the standard cross-entropy loss where the goal is to correctly classify nodes with the binary label of the QUBO solution $\bf x$}. For inference, we consider two modes to obtain an approximation of the solutions of QUBO problems: (1) Neural. As standard in neural networks, we employ QUBO-GNN to approximate the QUBO solution, and (2) Hybrid. We obtain a solution approximation using QUBO-GNN. This solution is considered an initial point in a classical solver. In our experiments, we used a small fixed number (up to 10) iterations of TS [18], as discussed in Section 3.1.
QUBO Solution Approximation.
We compare the approximated solution using various baselines in Table 1, where we consider small, medium, and large problems. As discussed in 2, we consider the case where the adjacency matrix $\bf A$} from Equation˜1 is constant, while the observed vector $\bf b$} is variable. To ensure statistical significance, for each size, we generate 10 instances of the matrix $\bf A$} and report the average and standard deviation metrics in Table 1 and Appendix C.2.
Our results indicate that: (i) Exhaustive search yields accurate solutions but is impractical for medium or large problems as the number of decision variables or nodes grows. (ii) Classical solvers achieve high accuracy on small problems, but their accuracy and efficiency decline as problem size increases. (iii) GNNs can generalize the QUBO problem solution by learning the map $q(\cdot)$, reducing inference time. Heterophily-designated GNNs like Dir-GNN and our QUBO-GNN offer higher accuracy and lower QUBO objectives. (iv) Combining neural approaches, such as our QUBO-GNN, with classical solvers like TS improves performance at a lower cost compared to classical solvers alone. Thus, our approach can serve as a strong prior for initializing classical solvers.
Table 2: The impact of adding the QUBO aware features from Equation˜14 to the node features.
QUBO Features Large $(k=1024)$
Acc(%) Rel. QUBO Inference
Objective Time (ms)
[…]
Figure 3: The relative QUBO loss convergence in training and validation of QUBO-GNN on the small problem ($k=10)$, with and without the QUBO-aware node features from Equation˜14.
QUBO Aware Features Improve Performance.
Our QUBO-GNN incorporates QUBO-aware node features in Equation˜14. In Table 2 and in the Appendix, we examine the impact of QUBO aware features on the downstream performance. Including QUBO-aware features adds negligible computation time (less than a millisecond) while improving accuracy by an average of 1.7%, reducing the relative QUBO objective by a factor of  2. Furthermore, as we show in Figure˜3, a beneficial property of the QUBO-aware features is their ability to accelerate the training convergence of the network.
8 Conclusions
This paper introduces a novel approach that merges graph learning with optimization problem-solving, offering a new perspective on approximating combinatorial problem solutions using GNNs. We analyze QUBO problems, revealing their piecewise constant sensitivity and heterophilic solutions, which guide the design of QUBO-GNN to approximate input-solution mappings.
We propose a scalable mechanism to efficiently generate input-solution data pairs without solving the NP-complete QUBO problem from scratch. Numerical results demonstrate QUBO-GNN’s superior accuracy and computational efficiency compared to other GNN solutions and classical solvers.
9 Limitations and Open Challenges
Our experiments demonstrate the ability to train a network to solve the QUBO problem Equation˜1 for a fixed matrix $\bf A$} and a variable vector $\bf b$}. Our approach, QUBO-GNN, generalizes to new and unseen $\bf b$} given fixed $\bf A$}, a major step beyond Schuetz et al. [46], where a new network must be trained for every pair $({\bf A},{\bf b})$. This setting, with fixed $\bf A$} and varying $\bf b$}, arises frequently in practice, e.g., asset allocation in Finance [21] and computational biology [30]. At the same time, while our theoretical and empirical findings establish this connection between GNNs and solving complex combinatorial problems, several aspects remain open. In particular, we identify the following research directions: (i) Variable Adjacency Matrix $\bf A$}.Allowing $\bf A$} to vary will enable the modeling of different underlying graphs and problems, rather than only changes in $\bf b$}. This is a well-studied challenge in non-neural approaches [38] and remains open for neural methods; (ii) Constrained Problems. We focus on QUBO, an unconstrained subclass, but many practical problems involve constraints, e.g., knapsack [25], topology optimization [49], and protein design [50]. Extending GNN-based methods to handle such constraints is a key direction; and (iii) Nonquadratic Problems. Beyond QUBO, future work can target nonquadratic binary optimization. For instance, [35] proposed methods for protein design [1] and telecommunication optimization [43]. Extending neural architectures to this broader class would further expand our framework’s applicability.
Appendix A Datasets
@@
We create 3 datasets: small ($k=10$), medium ($k=256$), and large ($k=1024$). For the small dataset, we generate exact solutions using an exhaustive search, with 1 million examples. For the small and medium examples, we generate 1000 examples using the procedure described in Algorithm˜2. Since we consider problems where $\bf A$} is fixed but the observed-vector $\bf b$} is variable, we repeat the process of generating datasets 10 times, so that we have examples on several randomly generated matrices $\bf A$}. The results in Section 7 are averaged across the 10 datasets, following the common experimental settings in Xu et al. [57]. In all cases, we employ a random training/validation split of 80%/20%, respectively.
@@
For the medium and large datasets, we chose $\bf A$} to be the matrix that discretizes the 2D Laplacian on a lattice. Such matrices often appear in thermodynamics calculations [15]. The data is generated using algorithm 2. In order to make the problem non-trivial, we choose $\sigma=4$, which makes roughly $5\%$ of the entries in ${\bf x}_{o}$ change. We have used $1000$ Tabu search steps in step (4) of the algorithm to approximate $\bf x$}.
Appendix B Hyperparameters
@@
In our experiments we consider the following hyperparameters for all GNN methods, including ours: ${\rm{learning\textunderscore rate}}\in\{1e-5,1e-4,1e-3\}$, ${\rm{weight}\textunderscore decay}\in\{1e-5,1e-4,0\}$, ${\rm{dropout}}\in\{0,0.1,0.5\}$. We train each network for up to 200 epochs. Dropout layers, if used, are applied after each hidden layer. We used the Adam [26] optimizer in all experiments.
@@
Appendix C Additional Results and Discussions
C.1 Trivial Cases of QUBO Solutions.
The minimization of ${\bf x}^{\top}{\bf A}{\bf x}$ can result in a solution that is either (i) all zeros, if the entries of A are all positive, (ii) all ones, if the entries are all negative, or (iii) a mix of ones and zeros which tend to be heterophilic. This type of behavior is known as phase transitions. The extreme cases (where all the values in $A$ are uniquely signed) are easy and are usually solved via rule-based mechanisms because they yield the trivial solution (see Beasley [2]). However, the heterophily designated GNNs in our experiments are known to handle both homophilic and heterophilic datasets. More critically, the difficult cases, as in the examples in our experiments, are in the phase transition zones, where indeed, the solution is heterophilic, as known from experiments in the literature and as analyzed in our paper.
C.2 Standard Deviation of the Relative QUBO Loss
@@
Table 3: The Relative QUBO loss (lower is better) average and standard deviations over 10 experiments, using various naive exhaustive search methods, classical methods, GNNs, and our QUBO-GNN.
Size $\rightarrow$ Small ($k=10$) Medium ($k=256$) Large $(k=1024)$
Method $\downarrow$
@@
GIN $6.02\cdot 10^{-1}\pm 5.10\cdot 10^{-2}$ $4.19\cdot 10^{-1}\pm 5.32\cdot 10^{-2}$ $6.20\cdot 10^{-1}\pm 6.73\cdot 10^{-2}$
@@
QUBO-GNN $2.83\cdot 10^{-3}\pm 4.48\cdot 10^{-4}$ $5.31\cdot 10^{-3}\pm 5.92\cdot 10^{-4}$ $1.26\cdot 10^{-2}\pm 5.47\cdot 10^{-3}$
QUBO-GNN + TS $1.37\cdot 10^{-5}\pm 2.77\cdot 10^{-6}$ $21\cdot 10^{-5}\pm 3.81\cdot 10^{-6} 1.1$$10\cdot 10^{-4}\pm 7.09\cdot 10^{-6}$
C.3 The Importance of QUBO-Aware Features
We now provide additional results on the importance of including the QUBO-aware features, on small and medium-sized problems, in Tables 4 and 5, respectively. As discussed in Section˜7, the results indicate that including the QUBO-aware features improves accuracy and the relative QUBO objective.
Table 4: The impact of adding the QUBO aware features from Equation˜14 to the node features on small sized problems ($k=10$).
@@ C.3 The Importance of QUBO-Aware Features
Table 5: The impact of adding the QUBO aware features from Equation˜14 to the node features on medium sized problems ($k=256$).